If the demand function faced by a firm is:

Q = 90 – 2P

TC = 2 + 57Q – 8Q2 + Q3

Determine the level of output at which the firm maximizes the profit.

Determine the best level of output for the above question by the MR and MC approach.



1
Expert's answer
2021-05-07T10:50:11-0400

the maximizing level of of utput Q is when "MC = MR = P"

"MC = TC' = 57 - 16Q +3Q^2"

"P= 45 - 0.5Q"

"45 -0.5Q = 57 -16Q +3Q^2"

"3Q^2 - 15,5Q +12 =0"

"Q = 4.21"